# Two-proton emission as source of spin-entangled proton pairs

### v2

@@ Abstract
Introduction. Spin correlation is a direct manifestation of quantum entanglement [1, 2, 3]. Violation of the local-hidden-variable (LHV) bound in terms of Clauser-Horne-Shimony-Holt (CHSH) formulation has been certified in atomic and optical systems [4, 5, 6, 7, 8]. Nuclear systems offer a conceptually distinct arena, where correlated fermions emerge naturally with strong interactions in a finite many-body environment [9, 10, 11]. A pioneering experiment by Sakai Sakai et al. measured the spin correlation of two protons emitted in the reaction $^{2}\mathrm{H}(p,{}^{2}\mathrm{He})n$. They observed a clear violation of the Bell-CHSH inequality [9], demonstrating the generation of spin-entangled proton pairs.
@@
Three-body model. We describe the $2p$ decay of 16Ne with a time-dependent three-body model consisting of the 14O core and two valence protons [28, 29, 30]. The three-body Hamiltonian is given as [28]
@@
with $\mu=m_{p}m_{C}/(m_{p}+m_{C})$, where $m_{C}$ and $m_{p}$ are the core and proton masses, respectively. The proton-proton interaction reads
$v_{pp}(\bm{r}_{1},\bm{r}_{2})=v_{pp,\mathrm{vac}}(\bm{r}_{1},\bm{r}_{2})+v_{pp,\mathrm{add}}(\bm{r}_{1},\bm{r}_{2}),$ (7)
where $v_{pp,\mathrm{vac}}$ is determined so as to reproduce the proton-proton scattering length, while the additional surface-dependent term $v_{pp,\mathrm{add}}$ accounts for medium effects induced by the core nucleus. Parameters of these potentials are explained later with the calculated results.
@@
Results. To clarify how the spin correlation $S(\Phi)$ reflects the underlying decay dynamics, we compare (i) a democratic-emission case and (ii) a sequential-emission case. Then, we later discuss (iii) a symmetric case, where the initial state of (i) is modified so as to lack the diproton correlation.
@@
In (ii) the sequential-emission case, the central term in the core-proton Woods-Saxon potential, $V_{WS}(r_{i})$, is multiplied by $1.158$. Owing to this deeper $V_{WS}(r_{i})$, the nuclear part of $v_{pp}$ is correspondingly weakened by the factor $0.2$ to keep the same $2p$ energy, $Q_{2p}=1.4$ MeV. With this setting, a sequential emission is promoted [14, 15, 16]. The mean lifetime in this case is evaluated as $\tau=9.1\times 10^{-20}$ seconds. In Fig. 2 (ii), the initial state shows a weaker diproton correlation, as confirmed by the suppressed asymmetry in the opening-angle distribution.
@@
Stability of the spin correlation is demonstrated in Fig. 5 (b), which shows the time evolution of $S(\Phi)$ at $\Phi=\pi/4$. In both of cases (i) and (ii), their values of $S(\Phi)$ remain constant over long times, indicating that the spin-correlation pattern shown in Fig. 5 (a) is robust throughout the decaying process. At later times, the emitted protons reach sufficiently large distances, where the subsequent dynamics is dominated by the spin-independent Coulomb interaction. Consequently, we expect that the spin correlations are preserved and accessible even at macroscopic distances. In this sense, case (i) realizes a femtometer-scale process, in which two protons are delivered to macrosopically separated locations keeping their spin entanglement. Therefore, a measurement of $S(\Phi)$ downstream of the decay process, as realized in the experimental geometry of Ref. [9], should directly access the pattern shown in Fig. 5 (a).
@@
Summary. We investigated the spin correlation in the $2p$ emission from 16Ne with a time-dependent three-body model, focusing on its sensitivity to the initial diproton correlation. When the decay proceeds as a democratic three-body process from an initial state containing a localized $S_{12}=0$ configuration, the spin correlation $S(\Phi)$ exhibits a characteristic angular dependence closely resembling that of a pure $S_{12}=0$ state. In contrast, when the decay is dominated by a sequential emission, the admixture of $S_{12}=1$ components washes out this pattern. Moreover, in the absence of an initial diproton correlation, a pronounced spin-correlation structure does not develop, even if the sequential decay is suppressed. As conclusion, democratic $2p$ emitters can serve as a source of spin-correlated proton pairs, and their spin correlation reflects the initial diproton correlation.
@@
Acknowledgments. We thank Hideyuki Sakai, Masaki Sasano, Simin Wang, and Tokurou Fukui for discussions. Following programs are acknowledged: (i) Multi-disciplinary Cooperative Research Program (MCRP) in FY2024 and FY2025 by Center for Computational Sciences, University of Tsukuba (project ID wo23i034); (ii) cooperative project of Yukawa-21 in Yukawa Institute for Theoretical Physics, Kyoto University.
Data availability. The data supporting the findings in this work are not publicly available. Those will be available from the authors upon reasonable request.

### v3

@@ Abstract
keywords:
Entanglement, Two-proton emission, Time-dependent calculation
journal: Nuclear Physics B
[…]
1 Introduction
Spin correlation is a direct manifestation of quantum entanglement [1, 2, 3]. Violation of the local-hidden-variable (LHV) bound in terms of Clauser-Horne-Shimony-Holt (CHSH) formulation has been certified in atomic and optical systems [4, 5, 6, 7, 8]. Nuclear systems offer a conceptually distinct arena, where correlated fermions emerge naturally with strong interactions in a finite many-body environment [9, 10, 11]. A pioneering experiment by Sakai Sakai et al. measured the spin correlation of two protons emitted in the reaction $^{2}\mathrm{H}(p,{}^{2}\mathrm{He})n$. They observed a clear violation of the Bell-CHSH inequality [9], demonstrating the generation of spin-entangled proton pairs.
@@ 1 Introduction
TABLE 1: Parameters for potentials. Note that the democratic and symmetric cases have the same set of parameters.
Democratic & Sequential
Symmetric
$V_{WS}$ $r_{0}$ [fm] $1.23$ $1.23$
$a_{0}$ [fm] $0.82$ $0.82$
$V_{0}$ [MeV] 42.25-42.25 48.93-48.93
[…]
$f_{pp}$ $1.0$ $0.2$
2 Three-Body Model
We describe the $2p$ decay of 16Ne with a time-dependent three-body model consisting of the 14O core and two valence protons [28, 29, 30]. The first-excited state of 14O is at $5.17$ MeV [35], which is higher than the experimental $2p$-emission energy of 16Ne, $Q_{2p}=1.4$ MeV [36, 37, 38, 39]. Thus, the effect of core excitation is expected as small, and neglected in this Letter.
The three-body Hamiltonian is given as [28]
@@ 2 Three-Body Model
with $\mu=m_{p}m_{C}/(m_{p}+m_{C})$, where $m_{C}$ and $m_{p}$ are the core and proton masses, respectively. These potentials are determined as
$V_{WS}(r)=\left[V_{0}+U_{ls}(\bm{l}\cdot\bm{s})\frac{1}{r}\frac{d}{dr}\right]\left(\frac{1}{1+e^{(r-R_{0})/a_{0}}}\right),$
$V_{Coul}(r)=\left\{\begin{array}[]{ll}\displaystyle\alpha\hbar c\frac{Z}{r}&(r\geq R_{0})\\ \displaystyle\alpha\hbar c\frac{Z}{2R_{0}}\left[3-\left(\frac{r}{R_{0}}\right)^{2}\right]&(r<R_{0})\end{array}\right$., (8)
where $R_{0}=r_{0}\cdot 14^{1/3}$, $\alpha=\frac{e^{2}}{4\pi\epsilon_{0}\hbar c}=\frac{1}{137.036}$, and $Z=8$. Other parameters are summarized in TABLE 1. On the other side, the proton-proton interaction includes the vacuum and additional terms [28]:
$v_{pp}(\bm{r}_{1},\bm{r}_{2})=v_{pp,\mathrm{vac}}(\bm{r}_{1},\bm{r}_{2})+v_{pp,\mathrm{add}}(\bm{r}_{1},\bm{r}_{2}).$ (9)
These terms read
$v_{pp,{\rm vac}}(\bm{r}_{1},\bm{r}_{2})=f_{pp}\cdot\Bigl[V_{R}e^{-a_{R}d^{2}}+V_{S}e^{-a_{S}d^{2}}\hat{P}_{S=0}\Bigr$.
$~~~~~~~~~~~~~~~~~~~~~~~~~~\Bigl.+V_{T}e^{-a_{T}d^{2}}\hat{P}_{S=1}\Bigr]+\frac{\alpha\hbar c}{d},$ (10)
$v_{pp,{\rm add}}(\bm{r}_{1},\bm{r}_{2})=f_{pp}\cdot w_{0}e^{-(R-R_{0})^{2}/B_{0}^{2}}\delta(\bm{r}_{1}-\bm{r}_{2}),$ (11)
where $d=\left|\bm{r}_{2}-\bm{r}_{1}\right|$ and $R=\left|(\bm{r}_{1}+\bm{r}_{2})/2\right|$. The operators $\hat{P}_{S=0}$ and $\hat{P}_{S=1}$ are the projectors to the spin-singlet and spin-triplet channels, respectively. Parameters for $v_{pp,{\rm vac}}$ were given as $V_{R}=200$ MeV, $V_{S}=-91.85$ MeV, $V_{T}=-178$ MeV, $a_{R}=1.487$ fm-2, $a_{S}=0.465$ fm-2, and $a_{T}=0.639$ fm-2, in order to reproduce the experimental scattering data of two protons [40]. For the additional term, $R_{0}=r_{0}\cdot 14^{1/3}$ and $B_{0}=0.6R_{0}$. This term accounts for medium effects to reproduce the experimental three-body energy, $Q_{2p}=1.4$ MeV, of 16Ne [36, 37, 38, 39]. Notice that $v_{pp,{\rm add}}(\bm{r}_{1},\bm{r}_{2})=0$ when $d\longrightarrow\infty$ or $R\longrightarrow\infty$. Thus, it does not affect the properties of two-body subsystems in vacuum. We also use the tuning factor $f_{pp}$ for the nuclear-force parts: see TABLE 1. Reason for employing these parameters is explained later with the calculated results and discussions.
@@
3 Results and Discussions
To clarify how the spin correlation $S(\Phi)$ reflects the underlying decay dynamics, we compare (i) a democratic-emission case and (ii) a sequential-emission case. Then, we later discuss (iii) a symmetric case, where the initial state of (i) is modified so as to lack the diproton correlation.
@@ 3 Results and Discussions
In (ii) the sequential-emission case, the central term in the core-proton Woods-Saxon potential is multiplied by $1.158$. Owing to this deeper $V_{WS}(r_{i})$, the nuclear part of $v_{pp}$ is correspondingly weakened by the factor $0.2$ to keep the same $2p$ energy, $Q_{2p}=1.4$ MeV: see TABLE 1. With this setting, a sequential emission is promoted [14, 15, 16]. The mean lifetime in this case is evaluated as $\tau=9.1\times 10^{-20}$ seconds. In Fig. 2 (ii), the initial state shows a weaker diproton correlation, as confirmed by the suppressed asymmetry in the opening-angle distribution.
@@
Stability of the spin correlation is demonstrated in Fig. 5 (b), which shows the time evolution of $S(\Phi)$ at $\Phi=\pi/4$. In both of cases (i) and (ii), their values of $S(\Phi)$ remain constant over long times, indicating that the spin-correlation pattern shown in Fig. 5 (a) is robust throughout the decaying process. At later times, the emitted protons reach sufficiently large distances, where the subsequent dynamics is dominated by the spin-independent Coulomb interaction. Consequently, we expect that the spin correlations are preserved and accessible even at macroscopic distances. In this sense, case (i) realizes a femtometer-scale process, in which two protons are delivered to macroscopically separated locations keeping their spin entanglement. Therefore, a measurement of $S(\Phi)$ downstream of the decay process, as realized in the experimental geometry of Ref. [9], should directly access the pattern shown in Fig. 5 (a).
@@
4 Summary
We investigated the spin correlation in the $2p$ emission from 16Ne with a time-dependent three-body model, focusing on its sensitivity to the initial diproton correlation. When the decay proceeds as a democratic three-body process from an initial state containing a localized $S_{12}=0$ configuration, the spin correlation $S(\Phi)$ exhibits a characteristic angular dependence closely resembling that of a pure $S_{12}=0$ state. In contrast, when the decay is dominated by a sequential emission, the admixture of $S_{12}=1$ components washes out this pattern. Moreover, in the absence of an initial diproton correlation, a pronounced spin-correlation structure does not develop, even if the sequential decay is suppressed. As conclusion, democratic $2p$ emitters can serve as a source of spin-correlated proton pairs, and their spin correlation reflects the initial diproton correlation.
The core nucleus 14O is treated as an inert core in this Letter. For investigating whether the core excitation disrupts the spin correlation or not, a technical improvement is necessary.
@@ 4 Summary
We thank Hideyuki Sakai, Masaki Sasano, Simin Wang, and Tokurou Fukui for discussions. Following programs are acknowledged: (i) Multi-disciplinary Cooperative Research Program (MCRP) in FY2024 and FY2025 by Center for Computational Sciences, University of Tsukuba (project ID wo23i034); (ii) cooperative project of Yukawa-21 in Yukawa Institute for Theoretical Physics, Kyoto University. The data supporting the findings in this work are not publicly available. Those will be available from the authors upon reasonable request.